OpenStudy (anonymous):

Does a=-1, b=-24, and c=144 of this quadratic equation?

OpenStudy (anonymous):

\[-y ^{2}=-24y+144\]

OpenStudy (anonymous):

y^2-24y+144=0 (y-12)(y-12)=0 y=12,12

OpenStudy (anonymous):

y2-24y+144 so YES!

OpenStudy (anonymous):

ax2+bx+c

OpenStudy (anonymous):

a=1

OpenStudy (anonymous):

I have to put what a, b and c are but i wasn't sure if any of these change to positive or if it is just how it looks.

OpenStudy (anonymous):

they need to be on the same side with 0 on the other

OpenStudy (anonymous):

so NO...either side, some are off

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

THanks! I haven't solved it yet but it looks like someone up there got 12, maybe that's the only repeating answer?

OpenStudy (anonymous):

yes answer is 12 if you solve it